# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

### v4

@@ 4.4.3. Qualitative Results
AMROD | FT of 0.9 | FT of 0.8 | FT of 0.7 | SR | DR | Mean | |
1 | 27.7 | ||||||
2 | 28.2 | ||||||
3 | 28.5 | ||||||
4 | 28.7 | ||||||
5 | 28.3 | ||||||
6 | 29.2 |
4.4.4. Ablation Study
@@
We conduct an ablation study to evaluate the impact of the OCL, AM, and ARR. As shown in Table 5, integrating the mean-teacher architecture achieves the performance of 25.9 mAP. Building upon this foundation, the proposed OCL, AM, and ARR achieve an additional 1.5, 1.4, and 0.8 mAP improvement individually. Moreover, the combination of the OCL and AM modules raises the performance to 28.6 mAP. Furthermore, the inclusion of the ARR mechanism brings a 0.6 mAP boost in performance, culminating in a comprehensive model performance of 29.2 mAP.
As presented in Table 6, we further perform a fine-grained comparison study to assess the effectiveness of the AM and ARR modules. This is done by replacing AM and ARR with a Fixed threshold (FT), stochastic restoration (SR) (Wang et al., 2022), or data-driven restoration (DR) (Brahma and Rai, 2023). The findings indicate that the dynamic nature of the AM module improves performance by 1.5, 1.0, and 0.7 mAP over FT set at 0.9, 0.8, and 0.7, respectively. While the randomness of SR and the gradient-only approach in DR limit the performance to 28.7 and 28.3 mAP, the ARR mechanism achieves the performance of 29.2 mAP. These results highlight the significance of the adaptability of AM and ARR over existing methods in dynamic environments. Additional ablation studies on hyper-parameters are provided in the appendix.
5. CONCLUSION
In this work, we proposed AMROD to address the two challenges in Continual Test-Time Adaptation (CTTA). Firstly, object-level contrastive learning leverages ROI features for contrastive learning to refine the feature representation, tailored for object detection. Secondly, the adaptive monitoring module enables efficient adaptation and high-quality pseudo-labels by dynamically skipping unstable adaptation and updating the category-specific threshold, based on the predicted confidence scores. Lastly, the adaptive randomized restoration mechanism selectively reset inactive parameters to mitigate forgetting while retaining valuable knowledge. The empirical results of four CTTA tasks for object detection demonstrate the efficiency and efficacy of AMROD in both short-term and long-term adaptation.
Appendix A Appendix Overview
We provide additional details and analysis in this appendix. In section B, we provide implementation details on AMROD and other baseline methodologies. In section C, we furnish complete results for the quantitative research, additional visualization examples, and additional ablation studies on hyper-parameters. In section D, we further discuss the continual learning. In section E, we discuss the limitations and prospective directions of our research.
Appendix B Additional Implementation Details
Following (VS et al., 2023b), we trained a Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) on the training set of Cityscapes (Cordts et al., 2016) as the source-trained model of Cityscapes. We utilized an SGD optimizer with a learning rate of 0.001, a momentum of 0.9, and a batch size of 1 and 70k iterations. Similarly, we trained another source model on the clear condition of the SHIFT (Sun et al., 2022) with 30 epochs. We will specify the implementation details of the AMROD and other baseline methods in the following subsections.
B.1. Implememtation of AMROD
Detailed hyper-parameters of AMROD for each benchmark are listed in Table 7. Furthermore, we present a pseudo-code outline of our AMROD framework in Algorithm 1. The three key components of AMROD are highlighted, i.e., Object-level Contrastive Learning, Adaptive Monitoring, and Adaptive Randomized Restoration.
B.2. Implememtation of Baseline Methods
The CoTTA (Wang et al., 2022) and SVDP (Yang et al., 2024) are originally designed for the classification or segmentation tasks. We implement them based on their public code, with some modifications to fit into the objection detection. Specifically, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a fixed threshold for pseudo-labeling and a stochastic restoration mechanism. Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on the mean-teacher framework. Further, the MemCLR (VS et al., 2023b) and IRG (VS et al., 2023a) are source-free domain adaptation methods for object detection. Therefore, we implement the MemCLR and IRG based on their public code, with little modifications to follow the CTTA setting (i.e., updating the model parameters during evaluation). For a fair comparison, we set the same momentum of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as AMROD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017). Other hyper-parameters are retained as specified in their public code.
Appendix C Additional Experiment Results
C.1. Quantitative Results
C.2. Qualitative Results
As shown in figure 4 - 8, we present more visualization results to compare the AMROD and other baseline methods based on the final round adaption on the long-term Cityscapes-to-Cityscapes-C task. The results further validate the effectiveness of AMROD in distinguishing more foreground object categories and better locating them, compared with other baseline methods.
C.3. Additional Ablation Study
To investigate the impact of hyperparameters, we conducted an ablation study on four key values on the long-term Cityscapes-to-Cityscapes-C task, including , , , and . The chosen hyper-parameter configuration in this paper of , , , and yielded a performance of 29.2 mAP. The ablation results, corresponding to different configurations, produced the following mAP scores: 28.8, 29.2, 29.0, 28.9, 29.1, 29.0, and 28.9, respectively. These findings highlight that the selected hyperparameters offer a balanced performance and confirm the stability of CTAOD, as it demonstrates minimal sensitivity to hyperparameter variations.
Hyper-parameter | Description | City2City-C-Short | City2City-C-Long | SHIFT-Short | City2ACDC-Long |
Learning rate | 0.001 | 0.001 | 0.0001 | 0.001 | |
Momentum in EMA | 0.9996 | 0.9998 | 0.9999 | 0.9998 | |
Temperature in CL | 0.07 | 0.07 | 0.07 | 0.07 | |
Initial valuse in DT | 0.8 | 0.8 | 0.8 | 0.8 | |
Hyper-parameter in DT | 0.95 | 0.95 | 0.95 | 0.95 | |
Hyper-parameter in DT | 1.3 | 1.3 | 1.3 | 1.3 | |
The upper-bound in DT | 0.9 | 0.9 | 0.9 | 0.9 | |
The lower-bound in DT | 0.7 | 0.7 | 0.7 | 0.7 | |
Hyper-parameter in DS | 0.7 | 0.75 | 0.6 | 0.7 | |
The threshold in DS | 1.5 | 1.4 | 1.4 | 1.4 | |
Quantile value | 0.01 | 0.001 | 0.001 | 0.01 |
Time | |||||||||||||||||||||||||||
Round | 1 | 2 | 3 | 4 | 5 | / | |||||||||||||||||||||
Condition | F. | M. | S. | B. | D. | F. | M. | S. | B. | D. | F. | M. | S. | B. | D. | F. | M. | S. | B. | D. | F. | M. | S. | B. | D. | / | / |
Source | 36.1 | 8.1 | 0.2 | 31.0 | 6.7 | 36.1 | 8.1 | 0.2 | 31.0 | 6.7 | 36.1 | 8.1 | 0.2 | 31.0 | 6.7 | 36.1 | 8.1 | 0.2 | 31.0 | 6.7 | 36.1 | 8.1 | 0.2 | 31.0 | 6.7 | / | / |
CoTTA | 38.2 | 10.6 | 0.4 | 33.8 | 9.3 | 39.9 | 10.7 | 0.5 | 34.0 | 9.9 | 40.6 | 10.4 | 0.6 | 34.4 | 9.8 | 40.9 | 10.4 | 0.6 | 35.4 | 9.1 | 40.8 | 10.4 | 0.6 | 36.5 | 9.2 | / | / |
SVDP | 36.8 | 8.8 | 0.6 | 43.5 | 11.2 | 43.9 | 11.8 | 1.7 | 47.9 | 14.9 | 46.2 | 13.9 | 2.5 | 50.0 | 17.8 | 46.2 | 14.4 | 3.2 | 49.0 | 19.6 | 45.0 | 15.4 | 3.8 | 47.8 | 19.9 | / | / |
IRG | 37.9 | 9.0 | 0.7 | 44.3 | 12.2 | 46.1 | 12.9 | 2.0 | 49.2 | 16.6 | 47.6 | 15.2 | 3.0 | 49.5 | 19.5 | 45.2 | 17.3 | 4.6 | 48.6 | 21.4 | 45.6 | 17.7 | 6.0 | 46.7 | 21.9 | / | / |
MemCLR | 37.7 | 8.9 | 0.8 | 45.5 | 13.5 | 45.8 | 13.6 | 2.3 | 49.8 | 17.7 | 47.3 | 16.4 | 4.1 | 50.2 | 20.3 | 46.4 | 17.4 | 4.8 | 47.8 | 21.9 | 45.0 | 18.1 | 5.1 | 46.8 | 22.4 | / | / |
Ours-unstop | 39.0 | 10.4 | 0.8 | 48.0 | 13.8 | 48.2 | 13.7 | 3.1 | 52.9 | 18.1 | 49.3 | 15.9 | 5.3 | 53.3 | 21.3 | 48.5 | 17.1 | 7.1 | 52.2 | 23.5 | 49.4 | 18.6 | 7.7 | 51.7 | 25.0 | / | / |
Ours | 39.2 | 9.9 | 0.8 | 46.7 | 12.7 | 46.8 | 13.5 | 2.7 | 52.3 | 17.9 | 49.1 | 15.8 | 4.5 | 53.9 | 21.2 | 48.8 | 17.7 | 6.9 | 53.8 | 23.1 | 49.4 | 18.5 | 8.0 | 52.9 | 24.7 | / | / |
Round | 6 | 7 | 8 | 9 | 10 | Mean | Gain | ||||||||||||||||||||
Source | 36.1 | 8.1 | 0.2 | 31.0 | 6.7 | 36.1 | 8.1 | 0.2 | 31.0 | 6.7 | 36.1 | 8.1 | 0.2 | 31 | 6.7 | 36.1 | 8.1 | 0.2 | 31.0 | 6.7 | 36.1 | 8.1 | 0.2 | 31.0 | 6.7 | 16.4 | / |
CoTTA | 39.6 | 10.3 | 0.6 | 34.9 | 9.4 | 40.2 | 10.8 | 0.5 | 35.3 | 10.0 | 40.0 | 10.8 | 0.5 | 35 | 9.7 | 40.1 | 10.5 | 0.5 | 34.1 | 9.5 | 40.8 | 10.4 | 0.5 | 36.3 | 9.6 | 19.1 | +2.7 |
SVDP | 44.4 | 15.6 | 4.2 | 47.6 | 20.5 | 44.1 | 16.1 | 4.3 | 46.8 | 20.4 | 43.0 | 15.9 | 4.6 | 45.7 | 20.7 | 42.6 | 16.0 | 4.8 | 45.0 | 20.2 | 41.5 | 16.2 | 4.6 | 44.6 | 20.2 | 25.3 | +8.9 |
IRG | 43.9 | 18 | 5.9 | 44.8 | 22.1 | 42.3 | 17.5 | 6.6 | 42.9 | 22.2 | 41.7 | 16.9 | 7.2 | 42.1 | 21.8 | 38.8 | 17 | 7.2 | 40.1 | 20.6 | 37.0 | 16.7 | 7.1 | 38.5 | 20.5 | 25.6 | +9.2 |
MemCLR | 43.4 | 19.3 | 5.9 | 44.6 | 23.3 | 41.1 | 19.5 | 6.6 | 43.1 | 23.9 | 39.6 | 19.1 | 7.0 | 40.9 | 24.3 | 38.6 | 18.9 | 7.2 | 39.1 | 22.4 | 36.8 | 18.5 | 7.5 | 38.4 | 21.8 | 26.0 | +9.6 |
Ours-unstop | 48.5 | 19.7 | 8.9 | 50.7 | 26.3 | 47.8 | 19.6 | 9.8 | 49.0 | 26.1 | 45.7 | 19.3 | 10.5 | 49.4 | 26.9 | 45.2 | 19.7 | 11.2 | 47.4 | 26.9 | 45.7 | 20.0 | 11.8 | 46.4 | 25.8 | 29.0 | +12.6 |
Ours | 48.6 | 19.8 | 9.1 | 51.2 | 25.4 | 48.8 | 20.1 | 10.0 | 49.5 | 27.0 | 47.3 | 19.9 | 10.0 | 49.2 | 27.9 | 46.9 | 19.4 | 11.0 | 47.5 | 28.9 | 46.2 | 19.5 | 11.3 | 47.3 | 29.1 | 29.2 | +12.8 |
Appendix D Additional Related Work
D.1. Continual Learning
Continual learning, also known as incremental learning or life-long learning, typically involves enabling the model to retain previously acquired knowledge while learning from a sequential series of tasks, i.e. preventing catastrophic forgetting (De Lange et al., 2021; Parisi et al., 2019). It is commonly categorized into replay methods (Rebuffi et al., 2017; Tiwari et al., 2022; Rolnick et al., 2019), parameter isolation method (Aljundi et al., 2017; Xu and Zhu, 2018), and regularization-based methods (Kirkpatrick et al., 2017; Li and Hoiem, 2017; Zenke et al., 2017). As an illustration, Elastic weight consolidation (EWC) (Kirkpatrick et al., 2017) is a regularization-based technique that penalizes the changing of parameters with a significant impact on prediction, based on the Fisher Information Matrix (FIM). In this paper, motivated by (Kirkpatrick et al., 2017; Brahma and Rai, 2023), we utilize gradients to approximate FIM as a metric of parameter importance for resetting noisy parameters. Furthermore, we introduce randomness to enhance model robustness during continual adaptation. Additionally, while the continual learning approaches aim to tackle catastrophic forgetting in sequences of new tasks, our work focuses on learning from different domains for a single task.
Appendix E Limitation
@@
Firstly, our experiments are primarily based on a single detector backbone, i.e. Faster R-CNN (Ren et al., 2015). Future work will extend to other advanced detector backbone networks to assess the universality and generalization capabilities of AMROD. Moreover, the proposed ARR mechanism to reset the model every iteration increases the computational cost. Future work could explore more efficient mechanisms for restoration, such as restoring the model only when significant changes in the distribution of the target domain are detected. Finally, the evaluation tasks in our work are designed to simulate real-world adaptation scenarios by incorporating datasets affected by corruption and adverse conditions. However, real-world data distributions are inherently more complex. Consequently, a promising avenue for future research would be to apply our methodology in practical, real-world systems to further validate its effectiveness.
Time | ||||||||||||||||||||||
Round | 1 | 2 | 3 | 4 | 5 | / | ||||||||||||||||
Condition | Fog | Night | Rain | Snow | Fog | Night | Rain | Snow | Fog | Night | Rain | Snow | Fog | Night | Rain | Snow | Fog | Night | Rain | Snow | / | / |
Source | 52.3 | 18.7 | 33.5 | 39.6 | 52.3 | 18.7 | 33.5 | 39.6 | 52.3 | 18.7 | 33.5 | 39.6 | 52.3 | 18.7 | 33.5 | 39.6 | 52.3 | 18.7 | 33.5 | 39.6 | / | / |
CoTTA | 53.7 | 19.7 | 38 | 42.4 | 53.6 | 20.3 | 37.7 | 42.2 | 53.7 | 20.1 | 38.1 | 43.3 | 53.1 | 19.7 | 37.7 | 42.9 | 52.6 | 19.5 | 37 | 42.2 | / | / |
SVDP | 52.8 | 20.0 | 35.6 | 42.0 | 53.4 | 22.0 | 37.4 | 43.8 | 54.0 | 23.0 | 38.8 | 44.0 | 54.6 | 23.5 | 38.7 | 43.8 | 53.4 | 24.0 | 38.0 | 44.4 | / | / |
IRG | 52.7 | 20.6 | 36 | 42.9 | 53.2 | 22.4 | 37.4 | 44.6 | 53.8 | 23.2 | 38.1 | 45.5 | 53.6 | 23.2 | 38.1 | 45.6 | 53.5 | 23.2 | 38.1 | 45.3 | / | / |
MemCLR | 52.9 | 21.2 | 35.2 | 42.8 | 52.5 | 23.1 | 37.3 | 44.3 | 53.9 | 23.2 | 38.4 | 44.8 | 52.8 | 23.2 | 38.2 | 44.4 | 51.5 | 23.5 | 37.2 | 43.8 | / | / |
Ours-unstop | 52.9 | 20.4 | 34.5 | 42.9 | 53 | 22.9 | 38.6 | 44.6 | 54.2 | 24 | 39.5 | 45.3 | 54.5 | 23.8 | 39.6 | 45.8 | 54.4 | 24.4 | 39.2 | 45.4 | / | / |
Ours | 52.8 | 20.1 | 35.2 | 42.9 | 53.6 | 22.2 | 38.3 | 44.2 | 53.9 | 23.5 | 39.6 | 45.0 | 54.4 | 24.6 | 38.6 | 45.5 | 54.6 | 25.1 | 39.2 | 46.3 | / | / |
Round | 6 | 7 | 8 | 9 | 10 | Mean | Gain | |||||||||||||||
Source | 52.3 | 18.7 | 33.5 | 39.6 | 52.3 | 18.7 | 33.5 | 39.6 | 52.3 | 18.7 | 33.5 | 39.6 | 52.3 | 18.7 | 33.5 | 39.6 | 52.3 | 18.7 | 33.5 | 39.6 | 36 | / |
CoTTA | 52.3 | 19.1 | 36.3 | 42 | 51.3 | 19 | 36.5 | 41.8 | 51.2 | 18.9 | 36.1 | 41.4 | 51.5 | 19.2 | 36.4 | 41.3 | 50.9 | 19.1 | 36 | 42.4 | 37.8 | +1.8 |
SVDP | 53.3 | 24.0 | 38.9 | 43.8 | 52.8 | 24.0 | 38.6 | 43.9 | 52.0 | 23.8 | 38.7 | 43.4 | 51.8 | 23.8 | 38.6 | 43.6 | 51.8 | 23.6 | 38.2 | 43.0 | 39.5 | +3.6 |
IRG | 52.6 | 23.5 | 37.5 | 43.7 | 51 | 23.1 | 37.1 | 42.8 | 50.2 | 23.1 | 37 | 42.2 | 49.8 | 23.2 | 36.5 | 41.1 | 49.7 | 22.5 | 36.5 | 40.7 | 38.9 | +2.9 |
MemCLR | 50.2 | 23.3 | 36.6 | 43 | 51.4 | 23.2 | 37.1 | 42.9 | 50.5 | 23.1 | 36.6 | 42.6 | 49.9 | 22.8 | 36.8 | 42 | 49.9 | 22.9 | 36.6 | 41 | 38.7 | +2.7 |
Ours-unstop | 53.9 | 24.9 | 39.2 | 45.1 | 53.5 | 25.1 | 39 | 45.3 | 53.3 | 24.6 | 39.3 | 45.3 | 53.2 | 24.5 | 39.3 | 44.9 | 52.5 | 24.3 | 38.8 | 44.7 | 40.3 | +4.3 |
Ours | 54.4 | 25.4 | 38.7 | 46.1 | 53.5 | 25.6 | 38.5 | 46.2 | 54.1 | 25.9 | 38.0 | 45.8 | 53.3 | 25.9 | 37.9 | 45.4 | 53.5 | 25.4 | 37.4 | 45.3 | 40.4 | +4.4 |

### v5

@@ V-C Qualitative Results
Mean-Teacher | OCL | AM | ARR | Mean | Gain |
25.9 | / | ||||
26.7 | +0.8 | ||||
27.3 | +1.4 | ||||
27.4 | +1.5 | ||||
28.6 | +2.7 | ||||
29.2 | +3.3 |
V-D Ablation Study
V-D1 Ablation Study Results on Model Components
@@
We conduct an ablation study to evaluate the impact of the OCL, AM, and ARR. As shown in Table VI, integrating the mean-teacher architecture achieves the performance of 25.9 mAP. Building upon this foundation, the proposed OCL, AM, and ARR achieve an additional 1.5, 1.4, and 0.8 mAP improvement individually. Moreover, the combination of the OCL and AM modules raises the performance to 28.6 mAP. Furthermore, including ARR mechanism brings a 0.6 mAP boost in performance, culminating in a model performance of 29.2 mAP.
AMROD | Student | FT of 0.9 | FT of 0.8 | FT of 0.7 | SR | DR | Mean |
27.9 | |||||||
27.7 | |||||||
28.2 | |||||||
28.5 | |||||||
28.7 | |||||||
28.3 | |||||||
29.2 |
V-D2 Ablation Study Results on Module Variants
As presented in Table VII, we further perform a fine-grained comparison study to assess the effectiveness of the modules of AMROD. This is done by replacing the teacher predictions with the student prediction in the Mean-Teacher framework [30], AM with a Fixed threshold (FT) strategy, and ARR with stochastic restoration (SR) [24] or data-driven restoration (DR) [49] mechanism. The findings indicate a variant utilizing student model predictions achieved only 27.9 mAP. This performance dip is likely because the student model, being actively updated, is less stable than the teacher model, which is updated via the EMA strategy to mitigate error accumulation. Furthermore, replacing the AM module with FT strategies (FT of 0.9, 0.8, and 0.7) resulted in mAP scores of 27.7, 28.2, and 28.5, respectively. These are notably lower than AMROD’s 29.2 mAP, underscoring the benefit of AM’s adaptability to varying model confidence across diverse categories and domains, which FT inherently lack. Similarly, when the ARR module is substituted with SR or DR, performance drop to 28.7 mAP and 28.3 mAP, respectively. This suggests that SR’s randomness might discard valuable domain-specific knowledge, while DR could retain noisy parameters. In contrast, AMROD’s ARR, which selectively resets inactive parameters with higher probability while integrating a stochastic component, better preserves essential knowledge and ensures robust adaptation. These results highlight the significance of the adaptability of AM and ARR over existing methods in dynamic environments.
V-D3 Ablation Study Results on Hyperparameter Impact
To investigate the impact of hyperparameters, we conduct an ablation study on six key hyperparameters on the long-term Cityscapes-to-Cityscapes-C task, revealing the model’s sensitivity and optimal configurations over the five target domains. These hyperparameters are varied while keeping others at their default settings, including , , , , , and . As shown in Figure 4, for the AM module, the dynamic threshold update rate (Figure 4a) shows peak performance at 0.95, suggesting that a balanced reliance on recent and historical category scores is beneficial. The initial threshold (Figure 4b) demonstrates strong robustness, while the linear projection factor (Figure 4c) is optimal at 1.3, indicating that a moderate scaling of predicted scores best guides threshold adaptation. For dynamic skipping, the update rate (Figure 4d) achieves its best at 0.75 and the stability range parameter (Figure 4e) performs best at 1.3, effectively balancing the reaction to mean score fluctuations for pausing adaptation. Finally, the threshold (Figure 4f) in ARR, which determines the q-quantile for resetting parameters, yields peak performance at 0.0001. Deviations from this value likely result in either insufficient resetting of inactive parameters or excessive resetting of potentially useful knowledge. These findings highlight that the selected hyperparameters offer a balanced performance and confirm the stability of AMROD, as they demonstrate minimal sensitivity to small hyperparameter variations. Full hyperparameter configurations are provided in the supplementary material.
VI Limitation
@@
Firstly, following prior works on TTA in objection detection [16, 43], our experiments are primarily based on a single detector backbone, i.e. Faster R-CNN [3] with ResNet50 [1]. Additionally, AMROD is specifically designed for object detection tasks and has demonstrated its effectiveness exclusively in this domain. Future work will extend to other advanced detector backbone networks or other vision tasks to assess the universality and generalization capabilities of AMROD. Moreover, the proposed ARR mechanism to reset the model every iteration increases the computational cost. Future work could explore more efficient mechanisms for restoration, such as restoring the model only when significant changes in the distribution of the target domain are detected. Finally, the evaluation tasks in our work are designed to simulate real-world adaptation scenarios by incorporating datasets affected by corruption and adverse conditions. However, real-world data distributions are inherently more complex. Consequently, a promising avenue for future research would be to apply our methodology in practical, real-world systems to further validate its effectiveness.
VII Conclusion
In this work, we propose AMROD to address the two challenges in Continual Test-Time Adaptation (CTTA). Firstly, object-level contrastive learning leverages ROI features for contrastive learning to refine the feature representation, tailored for object detection. Secondly, the adaptive monitoring module enables efficient adaptation and high-quality pseudo-labels by dynamically skipping unstable adaptation and updating the category-specific threshold, based on the predicted confidence scores. Lastly, the adaptive randomized restoration mechanism selectively reset inactive parameters to mitigate forgetting while retaining valuable knowledge. The empirical results of four CTTA tasks for object detection demonstrate the efficiency and efficacy of AMROD in both short-term and long-term adaptation.